Sir (du) p= du(due to temperature change)+F(v) where f(v) is change in u due to change in volume du p= nc v dt+f(v), nc p dt- nrdt =nc v dt+f(v) so c p -c v= R+f(v)/ndt From this i am not getting c p- c v= R for real gas.Please tell where i am wrong or is it that c p -c v is not equal to R for real gases

Sir  (du)p=du(due to temperature change)+F(v) where f(v) is change in u due to change in volume 
dup=ncvdt+f(v),  ncpdt- nrdt =ncvdt+f(v)
so cp-cv=R+f(v)/ndt From this i am not getting cp-cv=R for real gas.Please tell where i am wrong or is it that cp-cv is not equal to R for real gases

Grade:12th pass

1 Answers

Naveen Kumar
9 years ago
H= U+PV.............
dh=du+d(PV)........
(Here we are putting PV=nRT which is ideal gas equation and not true for real gases in general)
dh=du+d(RT)..........
dh=du+RdT.;........
frac{ dh }{ dT }=frac{ du }{ dT }+R
C p=C v+R..........
C p-Cv=R
.....................................................................................................